OpenStudy (anonymous):

Circle P has a radius of 8 units with center P at (2, -1). Which equation defines circle P? Answer (x – 2)2 + (y + 1)2 = 64 (x + 2)2 + (y – 1)2 = 64 (x – 2)2 + (y + 1)2 = 8 (x + 2)2 + (y – 1)2 = 8

OpenStudy (anonymous):

(x – 2)2 + (y + 1)2 = 64

OpenStudy (anonymous):

how did u fined that so i can do the next one by my self

myininaya (myininaya):

\[(x-h)^2+(y-k)^2=r^2 => (h,k)=center ; r=radius\]

OpenStudy (anonymous):

(x-a)^2 + (y-b)^2 = r^2 where (a,b) is center of circle and r = radius.

OpenStudy (anonymous):

The reason is as follows: